OpenStudy (anonymous):

The Waltons, a family of 12, are being given a $330 government tax deduction. They are given $30 for each female and $20 for each male. Build a system of equations that can be used to find the number of females and males in the Walton family.

OpenStudy (anonymous):

Let the number of females be f. Let the number of males be m. Then the tax = 330 = 30f+20m Also f+m = 12. Thus m = 12-f Substituting in above: \[330=30f+20(12-f)\] \[330=10f+240\] \[90=10f\] \[f=9\] and \[m=10\]

OpenStudy (anonymous):

Oh, my bad. m = 3

OpenStudy (anonymous):

there are only 12 in the family. Your answer shows 19, the answer is 3 males @ $60 and 9 females @ 270. This equals $330.

OpenStudy (anonymous):

Yes, my bad. See above comment